# Stochastic Periodic Solutions for Newtonian Systems via Lyapunov Function

### v1

@@ Abstract
This paper establishes an existence theory for distributed periodic solutions to Newton’s equation with stochastic time-periodic forcing, where the friction matrix is the Hessian of a twice continuously differentiable friction function. Employing the Khasminskii criterion and Lyapunov functions, we prove its existence under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and their gradient inner product grows at least like an even-power polynomial. Provided the potential grows sufficiently fast at infinity, this result persists under bounded friction matrix perturbations. This largely confirms the stochastic Levinson conjecture proposed in [9] and resolves the open problem and the interesting problem posed in [9, P. 342].
Distributed periodic solutions exist in two key cases:
@@
Keywords: Stochastic time-periodic Newtonian systems; Stochastic Levinson’s conjecture; Damping Hamiltonian systems; Periodic solutions in distribution; Lyapunov’s method
@@ 1 Introduction
$\ddot{x}+A(x,\dot{x},t)\dot{x}+\nabla V(x)=e(t),$ (1.1)
with friction must have a periodic solution. That is, the system (1.1) admits a periodic solution if the friction matrix $A(x,\dot{x},t)$ is symmetric and positive definite and $A(x,\dot{x},t),e(t)$ are periodic in time $t$. Mathematically in 1940s, Levinson [14] conjectured the principle is true, which brought important dissipativeness theory (see [15, 2, 23, 24] and references therein). For deterministic Newtonian systems, Li et al. [15] verified Levinson’s conjecture by [15, Theorem 4.2, P. 2619]:
@@
$A(x,\dot{x},t)\geq\alpha I_{n};\quad x^{\top}\nabla V(x)\geq\eta\|x\|^{2}\,\vee\,\gamma\,x^{\top}A(x,\dot{x},t)\,x,\qquad\hbox{for}~{}\,\|x\|\geq l,$ (1.2)
@@
Since then, the stochastic Levinson conjecture has been widely studied. Jiang and Li [8] established sufficient conditions for the conjecture to hold, assuming the potential function exhibits at most quadratic growth and that the friction matrix is independent of $x,\dot{x}$. Ji et al. [10] investigated the existence of periodic probability solutions for stochastic differential equations (SDEs) with irregular coefficients using time-periodic Fokker-Planck equations. Subsequently in [11], they demonstrated the convergence of these solutions for SDEs with less regular coefficients. As an application, both papers [10, 11] provided sufficient conditions for the stochastic Levinson conjecture to hold, leveraging the Lyapunov function constructed by Wu [22]. For the theories of more complex time-recurrence solutions in distribution, we can refer to [3, 19, 4, 7, 18] and their many references therein.
@@
$\ddot{x}+\gamma(t)\,\dot{x}+\nabla_{x}V(x,t)=\dot{B}_{t},$ (1.5)
@@
This paper aims to establish the existence of periodic solutions in distribution for stochastic time-periodic Newtonian systems with either relaxation oscillating friction, or uniformly positive friction, or beyond.
We first present a general existence criterion for distributed periodic solutions to stochastic time-periodic Hessian-driven friction Newtonian systems in which either the potential force or the friction exhibits at least polynomial growth at infinity:
@@
The system (1.6) admits at least a $\mathcal{T}$-periodic solution in distribution if it satisfies Hypotheses (H1)-(H5). Here
(H1) All coefficients in system (1.6) are continuous in their respective domains and locally Lipschitz continuous in $(x,y)$. Let $F\in C^{2}(\mathbb{R}^{n}),V(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})$, and $E(x,y,t)$, $\Sigma(x,y,t)$ also be periodic in time $t$ with period $\mathcal{T}$. Additionally, $V(x,t)$ is lower bounded.
@@
$\lim_{\|x\|\rightarrow+\infty}\Big{[}V(x,t)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}=+\infty,\qquad\hbox{uniformly in}~{}t\in\mathbb{R}^{+}.$
@@
$-\langle\nabla_{x}V(x,t),\nabla F(x)-ax\rangle\leq M-b\|x\|^{2m},\qquad~{}\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R}^{+}.$
@@
$|V_{t}(x,t)|\leq c_{1}\,b\|x\|^{2m}+M_{1},\qquad\quad\qquad\qquad\qquad~{}~{}\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R}^{+},$
$\|\Sigma\|_{2}^{2}:={\rm Tr}[(\Sigma\Sigma^{\top})(x,y,t)]\leq 2c_{2}\,(a\|y\|^{2}+b\|x\|^{2m})+M_{2},\quad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}.$
@@
Under this minimal assumption alone, we use Theorem 1.2 to prove the stochastic Levinson conjecture holds, which greatly improves upon the counterpart result in [15]:
@@
where $c$ is given in (3.10). If both $Q_{2q}(x)$ and $P_{2p}(x)$ are positive definite and either $Q_{2q}(x)=\|x\|^{2q}$ or $P_{2p}(x)=\|x\|^{2p}$, then system (1.6) admits a $\mathcal{T}$-periodic solution in distribution.
@@
Finally, we prove the existence of a periodic solution in distribution for (1.10) with a uniformly positive friction and a potential of degree exceeding $2$.
@@
The system (1.10) admits at least a $\mathcal{T}$-periodic solution in distribution if it satisfies the Assumptions (A1)-(A4). Here
(A1) All coefficients in system (1.10) are continuous in their respective domains and locally Lipschitz continuous in $(x,y)$. Let $V(x)\in C^{2}(\mathbb{R}^{n})$ be lower bounded and $C(x,y,t),E(x,y,t)$ $\Sigma(x,y,t)$ also be periodic with respect to time $t$ with period $\mathcal{T}$.
@@
$C^{s}(x,y,t)\geq 2\alpha I_{n};\ \ \|C(x,y,t)\|\,\vee\,|E(x,y,t)\|\leq\beta,\qquad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+},$
@@
$\|\Sigma\|_{2}^{2}(x,y,t)\leq c\alpha\,(\|y\|^{2}+b\|x\|^{2+\epsilon})+M_{1},\qquad\qquad~{}\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}.$
The proofs of Theorems 1.2-1.4 rely on Khasminskii’s criterion [12, Theorem 3.8] via a delicate Lyapunov function construction, building upon ideas from Jiang et al  [^1] On the exponential ergodicity and large deviations of stochastic Hessian-driven damping Hamiltonian systems. Preprint.. Our polynomial Lyapunov functions are smaller than the exponential functions introduced by Wu [22]. Polynomial functions can offer superior probabilistic properties: their slower growth reduces the possibility of exponential explosion and broadens the admissible class of initial distributions.
@@
    In contrast, our main Theorem 1.2 proves that distributed periodic solutions exist if $V(x,t)+aF(x)\to\infty$ uniformly in $t$ as $\|x\|\to\infty$ and the friction-potential gradient inner product grows at least like an even-power polynomial. The result holds even under bounded friction matrix perturbations (requiring higher potential growth). This result confirms the stochastic Levinson conjecture under broader conditions, including cases not covered previously:
@@
    Theorem 1.1 requires strong dissipativity: the second inequality in (1.2) implies that the potential grows at least quadratically at infinity and has a higher degree than the friction (i.e., $p>q$). Consequently, it applies only to strongly dissipative deterministic time-periodic Newtonian systems. This excludes seminal cases like the time-periodically forced van der Pol equation — a milestone in dynamical systems (see [25, 26, 27]). Furthermore, the first inequality in (1.2) mandates that friction admits a positive lower bound.
@@
    Moreover, Theorem 1.3 resolves Open Problem (1.4) posed in [9, Section 4].
@@
The organization of this paper is as follows. In Section 2, necessary preliminaries are given. In Section 3, we will give the proofs of Theorem 1.2-1.4. In Section 4, we will provide a series of applications originating from plasma physics. In Section 5, we summarize our main results and state our further perspectives on stochastic time-recurrent Newtonian systems.
@@ 2 Preliminaries
This section is devoted to provide preliminaries for our setup. Firstly, the definition of periodic solution in distribution is introduced, which is referred to [3, 8, 9, 12, 10, 11, 5, 16, 17]. Then, we present Khasminskii’s criterion in [12, Theorem 3.8], which provides simple conditions for the existence of periodic solutions of SDEs in terms of Lyapunov functions, which is generalized by Ji et al. [10, 11] .
@@
At first, we introduce the following notation. Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in\mathbb{R}^{+}},\mathbb{P})$ be a filtered probability space satisfying the usual conditions. For periodic stochastic systems, the period of function $f$ with respect to time $t$ being $\mathcal{T}$ means that $f(x,t+\mathcal{T})=f(x,t)$ holds for all $t\in\mathbb{R}^{+}$ and $x\in\mathbb{R}^{n}$. Denote by $C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R}^{+})$ the space of $\mathcal{T}$-periodic and continuous functions on $\mathbb{R}^{n}\times\mathbb{R}^{+}$ and by $C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})$ the space $C^{2,1}(\mathbb{R}^{n}\times\mathbb{R}^{+})\cap C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R}^{+})$.
Consider the following $\mathcal{T}$-periodic stochastic differential equation
$\begin{cases}dZ_{t}=b(Z_{t},t)\,dt+\sigma(Z_{t},t)\,dB_{t},~{}~{}\quad t>s,\\ Z_{s}=z,\end{cases}$ (2.1)
where the drift $b:\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}^{n}$ and noise intensity $\sigma:\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}^{n\times n}$ are $\mathcal{T}$-periodic in their second variable $t$. Let $Z^{s,z}(t)$ be the solution of (2.1) and ${\mathbb{P}}(s,z,t,A):=\mathbb{P}\big{(}Z^{s,z}(t)\in A\big{)}$ its transition probability function. Then ${\mathbb{P}}(s,z,t,A)$ is $\mathcal{T}$-periodic in the sense that ${\mathbb{P}}(s+\mathcal{T},z,t+\mathcal{T},\cdot)\equiv{\mathbb{P}}(s,z,t,\cdot)$ for all $z\in\mathbb{R}^{n}$ and $0\leq s\leq t$, see [12, Theorem 3.4]. We call a Markov process $Z(t)$ to be $\mathcal{T}$-periodic if its law ${\mathbb{P}}_{0}(t,A):={\mathbb{P}}\{Z(t)\in A\}$ satisfies the equation
${\mathbb{P}}_{0}(s,A)=\int_{\mathbb{R}^{n}}{\mathbb{P}}_{0}(s,dz)\,{\mathbb{P}}(s,z,s+\mathcal{T},A)\equiv{\mathbb{P}}_{0}(s+\mathcal{T},A),\quad\forall s\geq 0,\ A\in\mathcal{B}(\mathbb{R}^{n}),$ (2.2)
see [12, P. 64,  (3.17)]. (2.2) is equivalent to for any $s\geq 0$, the laws of $Z(s)$ and $Z(s+\mathcal{T})$ are identical. According to [8, 9], such a $\mathcal{T}$-periodic solution $Z(t)$ is called to be periodic solution in distribution or distributed periodic solution.
Now, we give the Khasminskii criterion in [12, Theorem 3.8], which implies the existence of $\mathcal{T}$-periodic solution of Eq. (2.1). Moreover, it plays a crucial role in the proof of Theorem 1.2.
@@
([12, Theorem 3.8]) Suppose that the coefficients of (2.1) are $\mathcal{T}$-periodic in time $t$ and satisfy the local Lipschitz conditions, and suppose further that there exists a non-negative function $\Psi(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n},\mathbb{R}^{+})$, which satisfies
$\lim_{R\to\infty}~{}\inf_{\|x\|>R,~{}t\in\mathbb{R}^{+}}\Psi(x,t)=+\infty,$ (2.3)

### v2

# Stochastic Periodic Solutions for Newtonian Systems via Lyapunov Method

@@ Abstract
This paper establishes an existence theory for periodic probability solutions to both time periodically and stochastically forced Newton’s equation, where the friction matrix is the Hessian of a twice continuously differentiable friction function. Employing the periodic probability solution theory established by Ji et al [10, 11] and the Lyapunov method, we prove its existence under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and their gradient inner product grows at least like an even-power polynomial. Provided the potential grows slightly faster at infinity, this result persists under bounded friction matrix perturbations. This largely confirms the stochastic Levinson conjecture proposed in [9] with much less restrictions on frictions and potentials and resolves the open problem and the interesting problem posed in [9, P. 342].
Periodic probability solutions exist in two key cases:
@@
Keywords: Stochastic time-periodic Newtonian systems; Stochastic Levinson’s conjecture; Damping Hamiltonian systems; Periodic probability solution; Lyapunov’s method
@@ 1 Introduction
$\ddot{x}+A(x,\dot{x},t)\dot{x}+\nabla V(x)=e(t)$ (1.1)
with friction must have a periodic solution. That is, the system (1.1) admits a periodic solution if the friction matrix $A(x,\dot{x},t)$ is symmetric, positive definite and $A(x,\dot{x},t),e(t)$ are periodic in time $t$. Mathematically in 1940s, Levinson [14] conjectured the principle is true, which brought important dissipativeness theory (see [15, 2, 23, 24] and references therein). For deterministic Newtonian systems, Li et al. [15] verified Levinson’s conjecture by [15, Theorem 4.2, P. 2619]:
@@
$A(x,\dot{x},t)\geq\alpha I_{n};\quad x^{\top}\nabla V(x)\geq\eta\|x\|^{2}\,\vee\,\gamma\,x^{\top}A(x,\dot{x},t)\,x\qquad\hbox{for}~\ \|x\|\geq l$ (1.2)
@@
Since then, the stochastic Levinson conjecture has been widely studied. Jiang and Li [8] established sufficient conditions for the conjecture to hold, assuming the potential function exhibits at most quadratic growth and that the friction matrix is independent of $x,\dot{x}$. Ji et al. [10] studied the long-term behavior in the sense of distribution for stochastic differential equations (SDEs) with irregular coefficients using time-periodic Fokker-Planck equations. Instead of searching periodic solutions in distribution of SDEs, they worked with the periodic probability solutions of the associated Fokker-Planck equation, and established the existence theory of a periodic probability solution, which is applicable to both SDEs with locally Lipschitz continuous coefficients and SDEs with Sobolev coefficients. Subsequently in [11], they demonstrated the convergence of these solutions for SDEs with less regular coefficients. As an application, both papers [10, 11] provided sufficient conditions for the stochastic Levinson conjecture to hold, leveraging the Lyapunov function constructed by Wu [22]. For the theories of more complex time-recurrence solutions in distribution, we can refer to [3, 19, 4, 7, 18] and their many references therein.
@@
$\ddot{x}+\gamma(t)\,\dot{x}+\nabla_{x}V(x,t)=\dot{B}_{t}$ (1.5)
@@
We point out that under local Lipschitz continuity of the SDEs’ coefficients, the concept of a periodic solution in distribution (Jiang et al. [9, 8]) is subsumed by the concept of a periodic probability solution (Ji et al. [10, 11]). We refer the reader to Section 2 for details. This motivates us to investigate the stochastic Levinson conjecture in the broader sense by establishing the existence of periodic probability solutions for stochastic time-periodic Newtonian systems with either relaxation oscillating friction, or uniformly positive friction, or beyond.
We first present a general existence criterion for periodic probability solutions to stochastic time-periodic Hessian-driven friction Newtonian systems in which either the potential force or the friction exhibits at least polynomial growth at infinity:
@@
The system (1.6) admits at least a $\mathcal{T}$-periodic probability solution if it satisfies Hypotheses (H1)-(H5). Here
(H1) All coefficients in system (1.6) are continuous in their respective domains and locally Lipschitz continuous in $(x,y)$. Let $F\in C^{2}(\mathbb{R}^{n}),V(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$, and $E(x,y,t)$, $\Sigma(x,y,t)$ also be periodic in time $t$ with period $\mathcal{T}$. Additionally, $V(x,t)$ is lower bounded.
@@
$\lim_{\|x\|\rightarrow+\infty}\Big{[}V(x,t)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}=+\infty,\qquad\hbox{uniformly in}~t\in\mathbb{R}.$
@@
$\langle\nabla_{x}V(x,t),\nabla F(x)-ax\rangle\leq M-b\|x\|^{2m},\qquad~\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R}.$
@@
$|V_{t}(x,t)|\leq c_{1}\,b\|x\|^{2m}+M_{1},\qquad\quad\qquad\qquad\qquad~~\forall(x,t)\in\mathbb{R}^{n}\times\mathbb{R},$
$\|\Sigma\|_{2}^{2}:={\rm Tr}[(\Sigma\Sigma^{\top})(x,y,t)]\leq 2c_{2}\,(a\|y\|^{2}+b\|x\|^{2m})+M_{2},\quad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}.$
@@
Under this minimal assumption alone, we use Theorem 1.2 to prove the stochastic Levinson conjecture holds for polynomials $F$ and $V$, which greatly improves upon the counterpart result in [15]:
@@
where $c$ is given in (3.10). If both $Q_{2q}(x)$ and $P_{2p}(x)$ are positive definite and either $Q_{2q}(x)=\|x\|^{2q}$ or $P_{2p}(x)=\|x\|^{2p}$, then system (1.6) admits a $\mathcal{T}$-periodic probability solution.
@@
Finally, we prove the existence of a periodic probability solution for (1.10) with a uniformly positive friction and a potential of degree exceeding $2$.
@@
The system (1.10) admits at least a $\mathcal{T}$-periodic probability solution if it satisfies the Assumptions (A1)-(A4). Here
[…]
@@
$C^{s}(x,y,t)\geq 2\alpha I_{n};\ \ \|C(x,y,t)\|\,\vee\,|E(x,y,t)\|\leq\beta,\qquad\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R},$
@@
$\|\Sigma\|_{2}^{2}(x,y,t)\leq c\alpha\,(\|y\|^{2}+b\|x\|^{2+\epsilon})+M_{1},\qquad\qquad~\forall(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}.$
We note that Theorems 1.2-1.4 still hold under weaker regularity of the diffusion coefficient $A=(a^{ij})\in C_{\mathcal{T}}\left(\mathbb{R},W_{{\rm loc}}^{1,p}(\mathbb{R}^{n})\right)$ with all others continuous. We refer the reader to Theorem 2.3 for details.
The proofs of Theorems 1.2-1.4 rely on the periodic probability solution theory established by Ji et al. [10, 11] via a delicate Lyapunov function construction, building upon ideas from Jiang et al  [^1] On the exponential ergodicity and large deviations of stochastic Hessian-driven damping Hamiltonian systems. Preprint.. Our polynomial Lyapunov functions are smaller than the exponential functions introduced by Wu [22]. Polynomial functions can offer superior probabilistic properties: their slower growth reduces the possibility of exponential explosion and broadens the admissible class of initial distributions.
@@
    In contrast, our main Theorem 1.2 proves that distributed periodic solutions exist if $V(x,t)+aF(x)\to\infty$ uniformly in $t$ as $\|x\|\to\infty$ and the friction-potential gradient inner product grows at least like an even-power polynomial. The result holds even under bounded friction matrix perturbations (requiring higher potential growth). This result substantially advances the stochastic Levinson conjecture, including cases not covered previously:
@@
    Theorem 1.1 requires strong dissipativity: the second inequality in (1.2) implies that the potential grows at least quadratically at infinity and has a higher degree than the friction (i.e., $p>q$). Consequently, it applies only to strongly dissipative deterministic time-periodic Newtonian systems. This excludes seminal cases like the time-periodically forced van der Pol equation — a milestone in dynamical systems (see [25, 26, 27]). Furthermore, the first inequality in (1.2) mandates that friction admits a positive lower bound at a neighborhood of infinity.
@@
    In particular, Theorem 1.3 resolves Open Problem (1.4) posed in [9, Section 4].
@@
The organization of this paper is as follows. In Section 2, necessary preliminaries are given and discussions on relations of various distributed periodic solution concepts are provided. In Section 3, we will give the proofs of Theorems 1.2-1.4. In Section 4, we will provide a series of applications originating from plasma physics. In Section 5, we summarize our main results and state our further perspectives on stochastic time-recurrent Newtonian systems.
@@ 2 Preliminaries
This section is devoted to provide preliminaries for our setup. Firstly, the definition of periodic probability solution or periodic solution in distribution is introduced, which is referred to [3, 8, 9, 12, 10, 11, 5, 16, 17]. Then, we present [10, Theorem B], which provides conditions for the existence of periodic probability solutions of SDEs in terms of Lyapunov functions.
@@
At first, we introduce the following notation. Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\in\mathbb{R}},\mathbb{P})$ be a filtered probability space satisfying the usual conditions. For periodic stochastic systems, the period of function $f$ with respect to time $t$ being $\mathcal{T}$ means that $f(x,t+\mathcal{T})=f(x,t)$ holds for all $t\in\mathbb{R}$ and $x\in\mathbb{R}^{n}$. Denote by $C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R})$ the space of $\mathcal{T}$-periodic and continuous functions on $\mathbb{R}^{n}\times\mathbb{R}$ and by $C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$ the space $C^{2,1}(\mathbb{R}^{n}\times\mathbb{R})\cap C_{\mathcal{T}}(\mathbb{R}^{n}\times\mathbb{R})$. $C_{0}^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$ denotes the space of all those functions in $C^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$ with compact supports.
Ji et al [10, 11] considered the following $\mathcal{T}$-periodic stochastic differential equation
$\begin{cases}dZ_{t}=b(Z_{t},t)\,dt+\sigma(Z_{t},t)\,dB_{t},~~\quad t>s,\\ Z_{s}=z\end{cases}$ (2.1)
with Sobolev coefficients. They studied periodic solutions to (2.1) with irregular (in particular, non-Lipschitz) coefficients in the sense of distribution by its corresponding Fokker-Planck equation:
$\mathcal{L}^{*}p:=-\partial_{t}p+\partial_{ij}^{2}\left(a^{ij}p\right)-\partial_{i}(b^{i}p)=0,(x,t)\in\mathbb{R}^{n}\times\mathbb{R}.$ (2.2)
Here the diffusion matrix $(a^{ij})=\frac{1}{2}\sigma\sigma^{T}$ and the drift field $b$ are $\mathcal{T}$-periodic in its second variable, $\partial_{i}=\partial_{z_{i}},\partial_{ij}=\partial_{z_{i}z_{j}},i,j\in\{1,2,\cdots,n\}$, and the usual summation convention is used. The following definitions are adopted from [10, p.4-5].
Definition 2.1
(Periodic probability solution) A Borel measure $\mu$ on $\mathbb{R}^{n}\times\mathbb{R}$ is called a periodic probability solution to (2.2) if there is a family of Borel probability measures $\{\mu_{t}\}_{t\in\mathbb{R}}$ on $\mathbb{R}^{n}$ satisfying
$\mu_{t}=\mu_{t+\mathcal{T}},\ \forall t\in\mathbb{R},$
$a^{ij},b^{i}\in L_{{\rm loc}}^{1}\left(\mathbb{R}^{n}\times\mathbb{R},{\rm d}\mu_{t}{\rm d}t\right),\ \forall i,j\in\{1,2,\cdots,n\}$
and $\mathcal{L}^{*}\mu=0$ on $\mathbb{R}^{n}\times\mathbb{R}$ in the sense that
$\int_{\mathbb{R}}\int_{\mathbb{R}^{n}}\mathcal{L}\phi{\rm d}\mu_{t}{\rm d}t=0,\ \forall\phi\in C^{2,1}_{0}(\mathbb{R}^{n}\times\mathbb{R})$ (2.3)
such that ${\rm d}\mu={\rm d}\mu_{t}{\rm d}t$.
For a non-negative function $\Psi\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$ and each $\rho>0$, we define the $\rho$-sublevel set
$\Omega_{\rho}:=\{(z,t)\in\mathbb{R}^{n}\times\mathbb{R}:\Psi(z,t)<\rho\}.$
Definition 2.2
(Unbounded Lyapunov function) A non-negative function $\Psi\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n}\times\mathbb{R})$ is called an unbounded Lyapunov function with respect to $\mathcal{L}$ if there is a sequence $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$ of open sets in $\mathbb{R}^{n}$ satisfying $\mathcal{U}_{m}\subset\mathcal{U}_{m+1}\subset\subset\mathbb{R}^{n}$ for all $m\in\mathbb{N}$ and $\mathbb{R}^{n}=\bigcup_{m=1}^{+\infty}\mathcal{U}_{m}$ such that
${\rm inf}_{(\mathbb{R}^{n}\backslash\mathcal{U}_{m})\times\mathbb{R}}\Psi(z,t)\rightarrow+\infty,\ {\rm as}\ m\rightarrow+\infty,$ (2.4)
and there exist a $\rho_{k}>0$, called an essential lower bound of $\mathbb{R}^{n}$, and a constant $\gamma>0$, called a Lyapunov constant of $\Psi$, such that
$\mathcal{L}\Psi(z,t)\leq-\gamma,\ \forall(z,t)\in\left(\mathbb{R}^{n}\times\mathbb{R}\right)\backslash{\Omega_{\rho_{k}}}.$ (2.5)
The following is the existence theorem of periodic probability solutions to the Fokker-Planck equation (2.2) with Sobolev coefficients, due to [10].
@@
([10, Theorem B]) Let (H) hold. Suppose $A=(a^{ij})\in C_{\mathcal{T}}\left(\mathbb{R},W_{{\rm loc}}^{1,p}(\mathbb{R}^{n})\right)$ and $b\in C_{\mathcal{T}}\left(\mathbb{R}^{n}\times\mathbb{R}\right)$. If there is an unbounded Lyapunov function with respect to $\mathcal{L}$, then (2.2) admits a periodic probability solution. Here the hypothesis (H) is given as follows.
(H) Let $p>n+2$ and $a^{ij}\in L^{\infty}_{{\rm loc}}\left(\mathbb{R},W_{{\rm loc}}^{1,p}(\mathbb{R}^{n})\right)$ for each $i,j\in\{1,...,d\}$. The drift vector field $b=(b^{i})$ satisfies $b^{i}\in L^{p}_{\rm loc}\left(\mathbb{R}^{n}\times\mathbb{R}\right)$ for each $i\in\{1,...,d\}$.
Corollary 2.4
Suppose that the coefficients of (2.1) are $\mathcal{T}$-periodic in time $t$ and satisfy all regular assumptions in Theorem 2.3. If there exists a non-negative function $\Psi(x,t)\in C_{\mathcal{T}}^{2,1}(\mathbb{R}^{n},\mathbb{R})$, which satisfies
$\lim_{\|z\|\to\infty}~\Psi(z,t)=+\infty,\ {\rm uniformly\ in}\ t\in\mathbb{R}$ (2.6)